John is twenty years younger than Amy, and in five years' time he will be half her age. What is John's age now?
Found 2 solutions by ewatrrr, algebrahouse.com:
Answer by ewatrrr(24785)


Hi
John is twenty years younger than Amy
Let x and (x-20)represent Amy's and John's age NOW
Question states*** in 5 yrs...
(x-20) + 5 = (1/2)(x+5)
Solving for x
1/2x = 5/2 + 15
x = 5 + 30
x = 35, Amy's age now. John is now 15. (35-15)
CHECKING our Answer***in 5 yrs...
20 = (1/2)40

Answer by algebrahouse.com(1659)

"John is twenty years younger than Amy, and in five years' time he will be half her age. What is John's age now?"

x = Amy's age now
x - 20 = John's age now

x + 5 = Amy's age in 5 years
x - 15 = John's age in 5 years

x - 15 = 1/2(x + 5) {in five years, he will be half her age}
2x - 30 = x + 5 {multiplied entire equation by 2 to clear fraction}
2x = x + 35 {added 30 to both sides}
x = 35 {subtracted x from both sides}
x - 20 = 15 {substituted 35, in for x, into x - 20}

Amy is 35 now
John is 15 now
www.algebrahouse.com